OpenStudy (anonymous):

Simplify: (4a^5/3)^3/2

OpenStudy (anonymous):

sleeping again?

OpenStudy (saifoo.khan):

umm, idk..

OpenStudy (saifoo.khan):

pellet!

OpenStudy (anonymous):

\[(4a^{\frac{5}{3}})^{\frac{3}{2}}\]

OpenStudy (dumbcow):

\[(\frac{4a^{5}}{3})^{3/2} = \frac{(4a^{5})^{3/2}}{3^{3/2}} = \frac{4^{3/2}a^{5*3/2}}{3^{3/2}} = \frac{8\sqrt{a^{15}}}{\sqrt{27}}\]

OpenStudy (saifoo.khan):

still, is that a^(5/3) or whole thing over 3??

OpenStudy (anonymous):

(4a^5)^3 = 64a^15 3^3 = 27 So you have: 64a^15/sqrt(27). Now take the square root of this due to the fact that it is a exponent over 2. This gives you a final answer of 8*sqrt(a^15)/3 Hope this helps. If it does, please choose good answer. http://www.tutorsean.net

OpenStudy (anonymous):

\[4^{\frac{3}{2}}=\sqrt{4^2}=2^3=8\]

OpenStudy (anonymous):

and \[(a^{\frac{5}{3}})^{\frac{3}{2}}=a^{\frac{5}{2}}\]

OpenStudy (anonymous):

for a final answer of \[8a^{\frac{5}{2}}\]